OpenStudy (anonymous):

Solve the equation: (3y+2)(y+3)=y+14 Please show work

OpenStudy (anonymous):

multiply out the left side then make 1 side zero. solve the quadratic.

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

\[\Large (3y+2)(y+3)=y+14\] multiply left side and you'll get... \[\Large 3y^2+9y+2y+6=y+14\] from here you go... \[\Large 3y^2+11y+6-y-14=0\] we moved the left side to the right and changed signs... now you have quadratic equation... \[\Large 3y^2+10y-8=0\] Now use this formula... \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] you have a=3 b=10 and c=-8 just substitute and there you go !... ;)